Write the equation in standard form using only integers and a positive coefficient for x. Match your result to the correct answer below.
a..x + 3y = 6
b...x – 3y = 2
c...–x + 3y = –6
d...x – 3y = 6

Answer by Alan3354(69443)

Write the equation in standard form using only integers and a positive coefficient for x. Match your result to the correct answer below.
a..x + 3y = 6
b...x – 3y = 2
c...–x + 3y = –6
d...x – 3y = 6
----------------
If there was an equation given, it doesn't show.
---------------------
They all use integers.
The coeff of x in c is negative.
a, b and d match.